OpenStudy (anonymous):

y"-3y'+2y=0 y=0 and y'=2 when x=0

OpenStudy (anonymous):

What are the instructions it gives you?

OpenStudy (anonymous):

find the particular solution of the differential equation that satisfies the stated conditions...it's a exercise of the SWOKOWSKI book

OpenStudy (loser66):

it's a homogeneous , how to have particular solution which is just for non-homogenous form?

OpenStudy (loser66):

you have general solution only.

OpenStudy (whpalmer4):

Okay, assume you'll have a solution of the form \(y = e^{\lambda x}\) \[y'(x) = \lambda e^{\lambda x}\]\[y''(x) = \lambda^2 e^{\lambda x}\] \[y''(x) - 3y'(x) +2y(x) = 0\]\[\lambda^2 e^{\lambda x} - 3\lambda e^{\lambda x} + 2e^{\lambda x} = 0\]factoring you have \[e^{\lambda x}(\lambda^2 - 3\lambda + 2) = 0\]but \(e^{\lambda x} = y(x)\) so \[y(x)(\lambda^2 - 3\lambda + 2) = 0\]solutions to that are where \(y(x) = 0\) (nowhere) or roots of the quadratic, \(\lambda = 1, \lambda = 2\) So your general solution is \(y(x) = c_1 e^{\lambda x} + c_2 e^{\lambda x}\) or\[y(x) = c_1e^{x} + c_2e^{2x}\] Plug in your initial conditions to solve for the values of \(c_1, c_2\)

OpenStudy (whpalmer4):

\[y(0) = c_1e^{0} + c_2e^{2*0}\]\[y'(x) = c_1 e^x + 2c_2 e^{2x}\]\[y'(0) = 2 = c_1e^{0} + 2c_1 e^{2*0}\] That gives you enough information to solve for \(c_1,c_2\)

OpenStudy (whpalmer4):

isn't differentiation of \(e^{ax}\) great? :-)

OpenStudy (loser66):

oh, you don't know how to find characteristic equation, ok from the problem y'' stands for r^2 3y' stands for 3r 2y stands for 2 so, the characteristic equation is r^2 -3r +2 =0 solve for r , they are r1 = 1 and r2 =2 got me?

OpenStudy (loser66):

so that the general solution will have the form of \(C_1 e^{r_1x} + C_2e^{r_2x}\) which is \(y = C_1e^x +C_2e^{2x}\) so far not good???? hahahaha.... I am waiting for your confirm whether you got me or not

OpenStudy (loser66):

Ok, good. the sentence y =0, y' =2 when x =0 means y (0) = 0 y'(0) =2. Dat sit now, apply on our solution, we have y =..... that is the short form of y(x) = \(C_1e^x + C_2e^{2x}\) so, y (0)? just plug 0 into the solution where you see x so that y (0) =\(C_1 e^0 +C_2e^{2*0}= 0\) so, you have \(C_1 + C_2 =0 (let say (1)\)) take the derivative of the solution to get y'(x) = ...? and then plug 2 into it to get y'(2) =.... you will have another equation let say (2) solve (1) and (2) to get C1, C2 Plug them back to y (x) Done